A gas has a density of 3.84 g/L at a presuure of 3.50 atm and a temperature of 37 oC. What is the molar mass (g/mol) of this gas?

27.89

58.73

497.28

0.102
Answer #1

P = 3.5atm

T= 37.0 oC

= (37.0+273) K

= 310 K

density = 3.84 g/L

Lets derive the equation to be used

use:

p*V=n*R*T

p*V=(mass/molar mass)*R*T

p*molar mass=(mass/V)*R*T

p*molar mass=density*R*T

Put Values:

3.5 atm * MM = 3.84g/L * 0.08206 atm.L/mol.K *310.0 K

MM = 27.89 g/mol

Answer:27.89 g/mol